# A Multiobjective Mathematical Model for Optimal Irrigation Water Allocation

### v1

@@ 4 Formulations of the Mathematical Models and Solution Approach
$P_{c}$ Price of total crop production in Ton per hectare (T/ha)
@@
Hence, our consideration is to minimize the Environmental Flow Deficiency (EFD). Specifically, if the target environmental flow is set equal to or lower than the actual environmental flow, the EFD becomes zero. This restriction ensures that minimizing EFD is achieved, though it results in the minimum attainable net benefit.
@@
This multiobjective programming model simultaneously minimizes two conflicting goals: the negative Net Benefit function $f_{1}(X_{c},Env.Flow_{m})$ and the environmental flow deficiency function $f_{2}(X_{c},Env.Flow_{m})$. The constraints ensure feasible pumping limits, enforce both minimum and maximum land-use requirements, and guarantee that crop area and environmental flow allocations remain non-negative. In essence, the formulation captures the trade-off between maximizing agricultural returns and preserving ecological water requirements.
We develop code within the MATLAB environment and utilise MATLAB solvers to address the models. A comprehensive solution procedure and algorithm for solving Models 1–3 are detailed in Appendices 9.1 and 9.2.
@@ 5 Numerical Experiments, Results and Sensitivity Analysis
This section presents a numerical study that shows the optimal water allocation for irrigation can be achieved under a large number of constraints. The solution procedure given in (9.1) is used to maximize profit for Models (3) and minimize environmental deficiency of Model (4). MATLAB solvers are used, including ’fmincon’ with the interior-point method and the sequential quadratic programming (SQP) algorithm to solve the problems. For the multiobjective problem, we employ the MATLAB genetic algorithm and the Pareto search approach, along with the scalarization method described in [3]. These solvers are tested for their ability to handle large-scale optimization and the challenges of approximating the Pareto front. Among the tested options, we report the computational time taken and the approximation of the solution of the solvers used. All computations are carried out on an HP Evo laptop with $16$ GB RAM and a Core $i7$ processor running at $4.6$ GHz.
Case Study 5.1
This numerical example is based on the case study by Ullah and Nehring [28]. The study considers a regional water board managing ten main crops over 12 months. The total cultivable area is 23,076 hectares, with a maximum annual water pumping capacity of 50 GL and a minimum crop demand for each 1000 hectares. The costs for pumped water and surface water are 100,000 and 26,000 AUDAUD perper GLGL, respectively. The target environmental flow is set at 100 GL per month. Additional parameter data, including rainfall, evapotranspiration, crop coefficients, crop production, total crop income, and variable costs, are provided in [28, 29]. We are examining how to allocate crops across the available area while considering whether the environmental flow meets its target environmental flow requirement. We analyze cases both with and without satisfying this constraint to determine the highest possible profit. This analysis provides insights that can later be used to explore the trade-off between maximising profit and minimising environmental deficiency.
Table 1: Optimal Allocation of Crops and Environmental Flow for Model (3).
@@
Env. Jan Feb Mar April May June July Aug Sep Oct Nov Dec
Flow 0 0 0 0 0 0 0 0 0 0 0 0
We solved Model 3 and obtained the crop and environmental flow allocations shown in Table 1. In this case, the maximum profit achieved is $1.4968\times 10^{9}$. At this point, the maximum deficiency in environmental flow is $1,200$ GL.
Now we consider that the environment flow meets the target flow requirement of the environment in Model (3), that is,
@@
Table 2: Optimal crop and environmental flow allocation for Model (3) under constraint (6).
@@
Env. Jan Feb Mar April May June July Aug Sep Oct Nov Dec
Flow 100 100 100 100 100 100 100 100 100 100 100 100
We solved Model 3 using constraint (6) and obtained the crop and environmental flow allocations presented in Table 2. In this scenario, the maximum profit is $1.4080\times 10^{9}$. Here, the environmental flow deficiency is 0 GL, since constraint (6) does not allow for any shortfall.
We next minimize Model 4 with the same parameter settings as in Case Study 5.1, considering both the inclusion and exclusion of constraint (6). The optimal crop and environmental flow allocations, shown in Table 3, are identical in both situations. In this case, the total environmental flow deficiency is zero, and the profit obtained is $0.2742\times 10^{9}$. Note: The computational time taken to solve all three cases is approximately 8–10 seconds.
Table 3: Optimal crop and environmental flow allocation for Model (4).
@@
Env. Jan Feb Mar April May June July Aug Sep Oct Nov Dec
Flow 200 200 200 200 200 200 200 200 200 200 200 200
@@
  • Profit only: When the model focuses only on profit, nearly all available water is directed to the most valuable crops. This generates the highest income of $\mathdollar 1.4968\times 10^{9}$, but the river receives too little water, creating a total environmental flow shortage of $1,200$ GL.
@@ 6 Multiobjective Optimization Problems
Multiobjective optimization deals with finding the best balance between two or more objectives that often compete with each other. Because these goals conflict, there is rarely a single solution that is best for every objective at once. Instead, the aim is to identify a group of equally valid trade-off solutions called the Pareto front, where improving one objective would worsen at least one other. A common way to address such problems is scalarization, which reformulates a multiobjective problem as a single-objective one so that standard optimization techniques can be applied. In this work, we use the weighted-constraint method [3]. Here, the objective functions are expressed with adjustable weights: one weighted objective is optimized, while the remaining objectives are treated as constraints whose bounds depend on the chosen weights. By varying these weights evenly, we can approximate the complete Pareto-optimal set. We chose the weighted-constraint approach for its straightforward implementation and its ability to generate accurate Pareto solutions (more scalarization approaches in [23][4]). The detailed procedure for the two-objective case is described below.
@@
In Section 5, net profit and environmental deficiency are optimized separately, and the results highlight the trade-off between these two objectives. In Case Study 6.1, we solve the two-objective optimisation problem defined in (5).
Case Study 6.1
In this analysis, we use the input parameters from Case Study 5.1 to solve the multiobjective optimization problem defined in Model 3. Because it is a large-scale problem, different approaches are applied, including the ’gamultiobj’ and ’paretosearch’ solvers in the MATLAB environment, while the weighted-constraint method is employed to approximate the Pareto front.
We solve the problem described in (5) using the ‘gamultiobj‘ solver, increasing the population size up to $500$. The lower bounds are set as the minimum area for $X_{c}$ and zero for environmental flow, while the upper bounds are set at $5,000$ ha for $X_{c}$ and $300$ GL for environmental flow. The Pareto front obtained with this algorithm is shown in Figure 1(a), and the computation time along with the total number of Pareto points is reported in Table 4.
We now solve the problem defined in (5) using the ‘paretosearch‘ solver, with the population size increased to 500. The same lower and upper bounds are considered for $X_{c}$ and environmental flow values. The Pareto front approximated by this algorithm is shown in Figure 1(b), while the computation time and the total number of Pareto points are reported in Table 4.
Finally, we use the weighted-constraint scalarization approach. This method requires a sequence of steps detailed in Appendix 9.2. With weight normalisation as defined in $W$, the procedure involves solving subproblems (7) and (8) and comparing their solutions, discarding any non-dominated points. To solve the subproblems, we employ the fmincon solver with the ’Interior-Point’, ’SQP’, and ’SQP-Legacy’ algorithms, using the same lower and upper bounds as in the previous two cases. Following the algorithm outlined in Appendix 9.2, we then solve the problem defined in (5) to approximate the Pareto points, which are illustrated in Figure 1(c). A numerical summary is provided in Table 4.
@@
Comparison Analysis: We observe that the ’gamultiobj’ algorithm generates $345$ Pareto points in $15.72$ seconds under the same parameter settings. From Figure 1(a), the highest environmental deficiency level is around $130$ GL when profit is $1.91\times 10^{9}$, and the deficiency drops to zero when profit is between $(1.180\;\mbox{and}\;1.181)\times 10^{9}$. The ’paretosearch’ algorithm takes slightly more time but produces more Pareto points $500$ in total. In generating the Pareto front, what matters more than the number of points is how much of the Pareto front each algorithm can cover. For paretosearch, the highest deficiency is $1200$ GL at a profit of $1.49\times 10^{9}$, with deficiency reaching zero when profit is around $1.37\times 10^{9}$ (see, Figure 1(b)). Using the weighted-constraint scalarization approach, we find that the highest deficiency and corresponding profit are the same as ’paretosearch’. However, the deficiency starts to appear when the profit is $1.4\times 10^{9}$ and decreases to $0.2\times 10^{9}$ (see, Figure 1(c)).
To evaluate the true Pareto front, we refer to Case Study 5.1 results, where individual objectives $f_{1}$ and $f_{2}$ are optimised. The results show that the highest deficiency of $1200$ GL occurs at a profit of $1.49\times 10^{9}$, consistent with Figures 1(b) and (c) obtained for ’paretosearch’ and the ’scalarization’ approach, respectively. Moreover, maximising $f_{1}$ under target environmental flow constraints (6) in Case Study 5.1 results in zero deficiency when profit is $1.4\times 10^{9}$, which matches only the scalarization results in Figures 1(c). This indicates that ’paretosearch’ and ’gamultiobj’ fail to approximate the lower part of the Pareto front. Therefore, considering computation time, the number of generated points, and the accuracy of the true Pareto front, the scalarization approach performs better than the other methods.
Comparison analysis with the previous study [28]: The findings of the present study highlight several critical improvements over the earlier work of Ullah and Nehring [28]. A key point of divergence lies in the treatment of Environmental Flow Deficiency (EFD). Ullah and Nehring reported a maximum EFD of only 35 GL under a fixed monthly environmental flow target of $100$ GL. Given that EFD is defined as the shortfall between the prescribed environmental flow and the actual allocation, the theoretical maximum deficit for a uniform 100 GL monthly target is $1,200$ GL per year (see, obtained results in Table 1). This indicates that their model captured only a narrow portion of the feasible solution space. Consequently, the ecological shortfalls are understated, and the trade-offs between agricultural benefits and ecological requirements are not fully explored. In contrast, our scalarization-based approach succeeds in covering the entire feasible frontier, generating Pareto solutions that extend up to the theoretical maximum EFD. This broader coverage, as illustrated in Figure 1(c) (Pareto fronts generated by the Weighted-Constraint approach, $P_{w}^{k}$), provides a far more comprehensive representation of the trade-off surface compared to [28, Figure 1 in Page 10], where the solution set remains confined to a very limited region.
The two studies also differ markedly in terms of computational performance and solution diversity. Ullah and Nehring, using the NSGA-II evolutionary algorithm, reported only $34$ Pareto-optimal solutions, obtained at the expense of approximately $706.59$ minutes of computational time. In sharp contrast, the scalarization technique produces $989$ Pareto solutions in just $10.12$ seconds. This demonstrates not only substantial computational efficiency but also a significant improvement in the diversity of solutions available to decision-makers. The larger and more representative solution set provided by our approach ensures a more reliable understanding of trade-offs and expands the range of strategies that can be considered in irrigation planning.
Together, these results establish that the scalarization framework achieves superior performance in both accuracy and efficiency. By covering the full spectrum of environmental deficits and delivering a wide set of optimal solutions with minimal computational burden, our study offers a more rigorous and practical foundation for decision-making in the management of the Muhuri Irrigation Project.

### v2

@@ 5 Formulations of the Mathematical Models and Solution Approach
$P_{c}$ Price of total crop production in $AUD$ per hectare (AUD/ha)
@@
Model 2 is formulated as a single-objective optimization problem in which minimizing Environmental Flow Deficiency (EFD) is the sole objective. The target environmental flow $\text{Tar.Env.Flow}_{m}$ represents a minimum ecological requirement rather than an upper bound. Accordingly, the decision variable $\text{Env.Flow}_{m}$ may exceed this target when sufficient water is available or when environmental protection is prioritized. Any release above the target does not contribute to EFD, which is defined only as the shortfall below $\text{Tar.Env.Flow}_{m}$. Once the minimum environmental flow requirement is satisfied, EFD becomes zero for all feasible solutions.
For this reason, Model 2 is not intended to generate a realistic cropping plan in isolation. Instead, it serves as a boundary single-objective model that identifies the ecologically optimal extreme of the feasible solution space. In this formulation, crop allocation is governed by feasibility constraints such as minimum crop area requirements, total land availability, and pumping capacity rather than by economic incentives. The resulting solution is used as a reference point for evaluating trade-offs between economic and environmental objectives in the multiobjective formulation of Model 3.
@@
This multiobjective programming model simultaneously minimizes two conflicting goals: the negative Net Benefit function $f_{1}(X_{c},Env.Flow_{m})$ and the environmental flow deficiency function $f_{2}(X_{c},Env.Flow_{m})$. The constraints ensure feasible pumping limits, enforce both minimum and maximum land use requirements, and guarantee that crop area and environmental flow allocations remain non-negative. In essence, the formulation captures the trade-off between maximizing agricultural returns and preserving ecological water requirements.
We develop code within the MATLAB environment and utilise MATLAB solvers to address the models. A comprehensive solution procedure and algorithm for solving Models 1–3 are detailed in Appendices 10.1 and 10.2.
@@ 6 Numerical Experiments, Results and Sensitivity Analysis
This section presents a numerical study demonstrating that optimal irrigation water allocation can be achieved under a large number of operational constraints during dry year conditions. To account for the stochastic nature of the optimization solvers and their sensitivity to initial conditions, multiple randomized trials are performed. The solution procedure described in (10.1) is applied to solve the single objective formulations, where Model 1 (3), maximizes net economic benefit, and Model 2  (4), minimizes environmental flow deficiency. The single objective problems are solved using MATLAB nonlinear optimization solvers, including ‘fmincon’ implemented with the interior point method and the sequential quadratic programming (SQP) algorithm. For the multiobjective formulation, MATLAB’s genetic algorithm and Pareto search approach are employed, together with the scalarization method described in [3], to approximate the Pareto front and explore trade-offs between economic and environmental objectives. These solvers are evaluated based on their ability to handle large scale optimization problems and the challenges associated with approximating the Pareto front. All optimization experiments are conducted with randomly generated initial decision vectors and repeated multiple times to ensure robustness. For each solver and model combination, computational time and solution quality are recorded and summarized using representative statistics. All numerical computations are carried out on an HP Evo laptop equipped with $16$ GB RAM and a Core $i7$ processor operating at $4.6$ GHz.
Case Study 6.1
The numerical example is based on the framework developed by Ullah and Nehring [26], who applied an optimization model to the Muhuri Irrigation Project using dry-year data representative of low-flow, water-stressed conditions in Bangladesh. In the present study, dry-year scenarios are interpreted as operational worst-case conditions for irrigation planning, rather than as indicators of arid or climatically comparable environments. Accordingly, the modeling assumptions adopted here ensure methodological consistency and comparability with prior studies, while remaining aligned with the hydrological and institutional setting of the Muhuri Irrigation Project.
The analysis considers a regional water authority responsible for managing the ten most demand-intensive crops over a twelve-month planning horizon. The total cultivable area is 23,076 hectares, with a maximum annual groundwater pumping capacity of 50 GL and a minimum crop area requirement of 1,000 hectares per crop. The unit costs of pumped and surface water are set at 100,000 and 26,000 AUD per GL, respectively. The use of AUD, rather than Bangladeshi Taka (BDT) or USD, is maintained to ensure direct comparability with prior work. The target environmental flow is specified as 100 GL per month.(A fixed monthly environmental flow target is used to maintain consistency with the benchmark framework and to enable transparent comparison across alternative optimization formulations. Although environmental flow requirements in monsoon- dominated river systems are inherently seasonal, introducing seasonally varying targets would add hydrological complexity beyond the scope of this methodological study. Future work may incorporate seasonally adaptive environmental flow requirements.)
Key parameter values, including pumping limits and environmental flow targets, are adopted directly from existing studies [26, 27] to enable consistent benchmarking of alternative optimization formulations under identical operating conditions. These parameters are not intended as site-specific policy prescriptions, but rather as representative constraints for evaluating solution behavior and trade-off characteristics. All hydrological inputs, including river inflow, rainfall, and evapotranspiration, are adopted directly from the published datasets used in Ullah and Nehring [26], which were compiled from official project records of the Muhuri Irrigation Project. These data are treated as exogenous inputs to ensure methodological consistency and to focus the analysis on optimization behavior rather than hydrological estimation. Within this framework, allocation outcomes are examined both with and without enforcing the environmental flow constraint to identify the upper bound of attainable economic benefit and to explore trade-offs between profit maximization and environmental flow protection.
Table 1: Optimal Allocation of Crops and Environmental Flow for Model 1 (3).
@@
Month Jan Feb Mar April May June July Aug Sep Oct Nov Dec
E.Flow 0 0 0 0 0 0 0 0 0 0 0 0
EFD 100 100 100 100 100 100 100 100 100 100 100 100
We solved Model 1 (3) and obtained the crop and environmental flow allocations shown in Table 1. In this case, the maximum profit achieved is $\mathdollar 1.4968\times 10^{9}$. However, since the second objective of maintaining environmental flows was fully relaxed, no water was allocated to environmental requirements, and no pumping was required. As a result, the environmental flow deficit reaches its maximum possible value of 1200 GL
Now we consider that the environment flow meets the target flow requirement of the environment in Model 2 (4), and Model  3 (5), that is,
@@
Table 2: Optimal crop and environmental flow allocation for Model 2 (4) under constraint (6).
@@
Month Jan Feb Mar April May June July Aug Sep Oct Nov Dec
E.Flow 100 100 100 100 100 100 100 100 100 100 100 100
EFD 0 0 0 0 0 0 0 0 0 0 0 0
We solved Model 2 (4) using constraint (6) and obtained the crop and environmental flow allocations presented in Table 2. In this scenario, the maximum profit is $\mathdollar 1.4080\times 10^{9}$. Here, the environmental flow deficiency is 0 GL, since constraint (6) does not allow for any shortfall. To satisfy this requirement, additional water must be supplied through pumping when surface water is insufficient.
The crop area allocations reported in Tables 1 and 2 exhibit relatively tidy values, with several crops assigned areas close to the minimum allowable level. This outcome is a direct consequence of the modeling structure rather than an artefact of the solution process. In particular, the presence of minimum crop area constraints, together with upper bounds on individual crop areas and the absence of any objective term promoting crop diversification, leads the optimization algorithms to select boundary solutions. Once feasibility conditions are satisfied, continuous nonlinear solvers such as ‘fmincon’ typically converge to corner points of the feasible region. As a result, the reported allocations should be interpreted as constraint-driven extreme solutions that define the bounds of the feasible space, rather than as detailed operational cropping plans.
Table 3: Optimal crop and environmental flow allocation for Model 3 (5).
@@
Month Jan Feb Mar April May June July Aug Sep Oct Nov Dec
E.Flow 200 200 200 200 200 200 200 200 200 200 200 200
EFD 0 0 0 0 0 0 0 0 0 0 0 0
Next we minimize Model 3 (5) with the same parameter settings as in Case Study 6.1, considering both the inclusion and exclusion of constraint (6). The optimal crop and environmental flow allocations, shown in Table 3, are identical in both situations. In this case, the total environmental flow deficiency is zero, and the profit obtained is $\mathdollar 0.2742\times 10^{9}$. In Table 3, the environmental flow exceeds the target value of 100 GL per month and is uniformly allocated at 200 GL, reflecting surplus environmental releases under the multiobjective formulation rather than a change in the target environmental flow requirement. Note: The computational time taken to solve all three cases is approximately 8–10 seconds.
@@
  • Profit only: When the model focuses only on profit, nearly all available water is directed to the most valuable crops. This generates the highest income of $\mathdollar 1.4968\times 10^{9}$, but the river receives too little water, creating a total environmental flow shortage of $1200$ GL.
@@ 7 Multiobjective Optimization Problems
Multiobjective optimization deals with finding the best balance between two or more objectives that often compete with each other. Because these goals conflict, there is rarely a single solution that is best for every objective at once. Instead, the aim is to identify a group of equally valid trade-off solutions called the Pareto front, where improving one objective would worsen at least one other. A common way to address such problems is scalarization, which reformulates a multiobjective problem as a single-objective one so that standard optimization techniques can be applied. In this work, we use the weighted-constraint method [3]. Here, the objective functions are expressed with adjustable weights: one weighted objective is optimized, while the remaining objectives are treated as constraints whose bounds depend on the chosen weights. By varying these weights evenly, we can approximate the complete Pareto optimal set. We chose the weighted constraint approach for its straightforward implementation and its ability to generate accurate Pareto solutions (more scalarization approaches in [4], [22]). The detailed procedure for the two objective case is described below.
@@
In Section 6, net profit and environmental deficiency are optimized separately, and the results highlight the trade-off between these two objectives. In Case Study 7.1, we solve the two objective optimisation problem defined in (5).
Case Study 7.1
In this analysis, we use the input parameters from Case Study 6.1 to solve the multiobjective optimization problem defined in Model 3. Because it is a large scale problem, different approaches are applied, including the ’gamultiobj’ and ’paretosearch’ solvers in the MATLAB environment, while the weighted-constraint method is employed to approximate the Pareto front.
We solve the problem described in (5) using the ‘gamultiobj’ solver, increasing the population size up to $500$. The lower bounds are set as the minimum area for $X_{c}$ and zero for environmental flow, while the upper bounds are set at $5,000$ ha for $X_{c}$, (The per-crop upper bound is implemented as it ensures realistic crop distribution while preventing extreme monocropping and preserving comparability with the benchmark framework) and $300$ GL for environmental flow. The Pareto front obtained with this algorithm is shown in Figure 1(a), and the computation time along with the total number of Pareto points is reported in Table 4.
We now solve the problem defined in (5) using the ‘paretosearch’ solver, with the population size increased to 500. The same lower and upper bounds are considered for $X_{c}$ and environmental flow values. The Pareto front approximated by this algorithm is shown in Figure 1(b), while the computation time and the total number of Pareto points are reported in Table 4.
Finally, we use the weighted constraint scalarization approach. This method requires a sequence of steps detailed in Appendix 10.2. With weight normalisation as defined in $W$, the procedure involves solving subproblems (7) and (8) and comparing their solutions, discarding any non-dominated points. To solve the subproblems, we employ the fmincon solver with the ‘Interior Point’, ‘SQP’, and ‘SQP Legacy’ algorithms, using the same lower and upper bounds as in the previous two cases. Following the algorithm outlined in Appendix 10.2, we then solve the problem defined in (5) to approximate the Pareto points, which are illustrated in Figure 1(c). A numerical summary is provided in Table 4.
@@
Comparison Analysis: We observe that the ‘gamultiobj’ algorithm generates $345$ Pareto points in $15.72$ seconds under the same parameter settings. From Figure 1(a), the highest environmental deficiency level is around $130$ GL when profit is $\mathdollar 1.91\times 10^{9}$, and the deficiency drops to zero when profit is between ($\mathdollar 1.180\;\mbox{and}\;1.181)\times 10^{9}$. The ‘paretosearch’ algorithm takes slightly more time but produces more Pareto points $500$ in total. In generating the Pareto front, what matters more than the number of points is how much of the Pareto front each algorithm can cover. For paretosearch, the highest deficiency is $1200$ GL at a profit of $\mathdollar 1.49\times 10^{9}$, with deficiency reaching zero when profit is around $\mathdollar 1.37\times 10^{9}$ (see, Figure 1(b)). Using the weighted-constraint scalarization approach, we find that the highest deficiency and corresponding profit are the same as ’paretosearch’. However, the deficiency starts to appear when the profit is $\mathdollar 1.4\times 10^{9}$ and decreases to $\mathdollar 0.2\times 10^{9}$ (see, Figure 1(c)).
To evaluate the true Pareto front, we refer to Case Study 6.1 results, where individual objectives $f_{1}$ and $f_{2}$ are optimised. The results show that the highest deficiency of $1200$ GL occurs at a profit of $\mathdollar 1.49\times 10^{9}$, consistent with Figures 1(b) and (c) obtained for ‘paretosearch’ and the ‘scalarization’ approach, respectively. Moreover, maximising $f_{1}$ under target environmental flow constraints (6) in Case Study 5.1 results in zero deficiency when profit is $\mathdollar 1.4\times 10^{9}$, which matches only the scalarization results in Figures 1(c). This indicates that ‘paretosearch’ and ‘gamultiobj’ fail to approximate the lower part of the Pareto front. Therefore, considering computation time, the number of generated points, and the accuracy of the true Pareto front, the scalarization approach performs better than the other methods.
Comparison analysis with the previous study[26]: The findings of the present study highlight several critical improvements over the earlier work of Ullah and Nehring [26]. A key point of divergence lies in the treatment of Environmental Flow Deficiency (EFD). Ullah and Nehring reported a maximum EFD of only 35 GL under a fixed monthly environmental flow target of $100$ GL. Given that EFD is defined as the shortfall between the prescribed environmental flow and the actual allocation, the theoretical maximum deficit for a uniform 100 GL monthly target is $1200$ GL per year (see, obtained results in Table 1). This indicates that their model captured only a narrow portion of the feasible solution space. Consequently, the ecological shortfalls are understated, and the trade-offs between agricultural benefits and ecological requirements are not fully explored. In contrast, our scalarization based approach succeeds in covering the entire feasible frontier, generating Pareto solutions that extend up to the theoretical maximum EFD. This broader coverage, as illustrated in Figure 1(c) (Pareto fronts generated by the Weighted-Constraint approach, $P_{w}^{k}$), provides a far more comprehensive representation of the trade-off surface compared to [26, Figure 1 in Page 10], where the solution set remains confined to a very limited region.For cross method comparison, relative performance is discussed based on the shape and extent of the attainment surfaces rather than absolute magnitudes alone. The two studies also differ markedly in terms of computational performance and solution diversity. Ullah and Nehring, using the NSGA-II evolutionary algorithm, reported only $34$ Pareto optimal solutions, obtained at the expense of approximately $706.59$ minutes of computational time. In sharp contrast, the scalarization technique produces $989$ Pareto solutions in just $10.12$ seconds. This demonstrates not only substantial computational efficiency but also a significant improvement in the diversity of solutions available to decision makers. The larger and more representative solution set provided by our approach ensures a more reliable understanding of trade-offs and expands the range of strategies that can be considered in irrigation planning.
Together, these results establish that the scalarization framework achieves superior performance in both accuracy and efficiency. By covering the full spectrum of environmental deficits and delivering a wide set of optimal solutions with minimal computational burden, our study offers a more rigorous and practical foundation for decision making in the management of the Muhuri Irrigation Project.